A rectangular piece of sheet metal is 2 inches longer than it is wide. Squares measuring 4 inches on a side are cut from each corner and the sides are folded to form a box. If the volume of the box is 32 cubic inches, what are the dimensions of the sheet metal?
Answer by checkley71(8403)

THE HEIGHT OF THIS BOX IS 4 INCHES.
THE LENGTH IS W+2
THEREFORE THE VOLUME IS:
4*W(W+2)=32
4W^2+8W=32
4W^2+8W-32=0 FACTOR OUT 4
4(W^2+2W-8)=0
(W+4)(W-2)=0
W-2=0
W=2 ANSWER FOR THE WIDTH
L=2+2=4 FOR THE LENGTH.
PROOF
4*2*4=32
32=32